OpenStudy (anonymous):

A path for a new city park will connect the park enterance and Maine Street, as shown below. The path should be perpendicular to Maine Street. What is the equation that represents the path? y=________________

OpenStudy (anonymous):

this is for the problem stated above ^^^

OpenStudy (anonymous):

@SolomonZelman please help me

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

please note that the line "Main Street" passes at point (3,3) and (1,-1) so what is its slope?

OpenStudy (michele_laino):

hint: you have to apply this formula: slope=\[=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] where (x_1,y_1) and (y_1,y_2) are your points, namely (3,3) and (1,-1)

OpenStudy (anonymous):

2/1?

OpenStudy (michele_laino):

oops...where (x_1,y_1) and (x_2,y_2) are the points (3,3) and (1-1)

OpenStudy (michele_laino):

ok! I call with m the slope of "Main Street", so m=2

OpenStudy (michele_laino):

Now "Park entrance" has to be perpendicular with respect to "Main Street". If I call with m' the slope of the line "Park entrance" then we have this equation: \[m*m'=-1\] so, please compute m' from my equation

OpenStudy (michele_laino):

It is very simple, because I set m=2 into my formula and I get: \[2*m'=-1\] please compute m'

OpenStudy (anonymous):

ok so do you want me to substitute the new slope and sole for m?

OpenStudy (michele_laino):

ok! m'=...

OpenStudy (michele_laino):

if 2m'=-1, then \[m'=-\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

3

OpenStudy (michele_laino):

Now, in order to find the equation of "Park entrance", please you have to apply this equation: \[y-y _{1}=m'(x-x _{1})\] where m'=-1/2 and (x_1,y_1)=(0,4), since "park entrance" passes at point (0,4) as you can see from your drawing.So we have: \[y-4=-\frac{ 1 }{ 2 }x\] please, simplify that equation and you will get your answer